find length and width of area 204 perimeter 58
Answer by ankor@dixie-net.com(22740)

find length and width of area 204 perimeter 58
:
Let x = length, y = width
:
Perimeter:
2x + 2y = 58
Simplify, divide by 2:
x + y = 29
y = (29-x)
:
Area =
x*y = 204
;
Substitute (29-x) for y:
x(29-x) = 204
-x^2 + 29x - 204 = 0; a quadratic equation
:
Multiply by -1, (easier to factor):
x^2 - 29x + 204 = 0
:
Factor:
(x - 12)(x - 17) = 0
:
x = +12; then y = 17
or
x = 17; then y = 12